An association had a fund balance of 75 on 1/1 and 60 on 12/31. At the end of every month during this year, the association deposited 10. There were withdrawals of 5 on 2/28 and 25 on 6/30, 80 on 10/15 and 35 on 10/31. Find the dollar-weighted rate of return for the year.
Answer #1

This can be solve as follows

Using simple interest formula

75 + 10(1+(11/12)*i) + 10(1+(10/12)*i) + ... + 10(1+(1/12)*i) + 10 - 5(1+(10/12)*i) - 25(1+(6/12)*i) - 80(1+(2.5/12)*i) - 35(1+(2/12)*i) = 60

75+10*12 -5-25-80-35+(660/12)*i -(50/12)*i -(150/12)*i -(200/12)*i -(70/12)*i = 60

75+120-145+(660/12)*i -(470/12)*i = 60

(190/12)*i=10

i = 120/190 = 63.15%

This is the dollar weighted rate of return